OpenStudy (anonymous):

I am trying to find the derivative of an implicit equation of xe^y=x-y. After simplifying just the right side of the equation I get d/dx(xe^y)=1-dy/dx. A previous example shows this simplifying as d/dx(xe^y)=-dy/dx. What happened to the 1? I don't get it and am spending way too much time on this simple matter. Please help! John

OpenStudy (anonymous):

\[\Huge\bf \color{yellow}{Welcome~\to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~\to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~\to~\color{purple}{Open}}\color{blue}{Study!!!!}\]

OpenStudy (anonymous):

Do you mind posting a picture of your question?

OpenStudy (anonymous):

are you differentiating wrt 'x' then d/dx(xe^y)=1-dy/dx is correct now d/dx(xe^y) = e^y +x(e^y)*dy/dx ... (using product rule) so e^y +x(e^y)*dy/dx = 1-dy/dx now make dy/dx as the subject of

OpenStudy (anonymous):

Or taking a screenshot of you workings so far?

OpenStudy (turingtest):

indeed, the 1 should not be going anywhere in this case

OpenStudy (anonymous):

The book can also be wrong, I've seen this happen, you might want to check with your instructor then on that matter

OpenStudy (turingtest):

or seeing that 'previous example' you speak of may resolve the issue, as there may be a difference between the example and this problem that causes the 1 to drop out

OpenStudy (anonymous):

here is the example solved on this forum 2 years ago! ddx (xe y =x−y) ddx (xe y )=ddx (x−y) ddx (xe y )=ddx (x)−ddx (y) ddx (xe y )=dxdx −dydx ddx (xe y )=1−dydx ddx (xe y )=−dydx

OpenStudy (turingtest):

well then, I do think that's just wrong I'm afraid this site doesn't guarantee 100% accuracy, in exchange for free services I just it wasn't one I solved :P

OpenStudy (turingtest):

I just hope that that wasn't one I solved*

OpenStudy (anonymous):

Thanks for the replies from everyone. I am very new at this. John